# MCiteBench: A Benchmark for Multimodal Citation Text Generation in MLLMs

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have advanced in integrating diverse modalities but frequently suffer from hallucination. A promising solution to mitigate this issue is to generate text with citations, providing a transparent chain for verification. However, existing work primarily focuses on generating citations for text-only content, overlooking the challenges and opportunities of multimodal contexts. To address this gap, we introduce MCiteBench, the first benchmark designed to evaluate and analyze the multimodal citation text generation ability of MLLMs. Our benchmark comprises data derived from academic papers and review-rebuttal interactions, featuring diverse information sources and multimodal content. We comprehensively evaluate models from multiple dimensions, including citation quality, source reliability, and answer accuracy. Through extensive experiments, we observe that MLLMs struggle with multimodal citation text generation. We also conduct deep analyses of models’ performance, revealing that the bottleneck lies in attributing the correct sources rather than understanding the multimodal content.
@@
MCiteBench: A Benchmark for Multimodal Citation Text Generation in MLLMs
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have shown remarkable progress in integrating external information from diverse modalities, allowing them to generate responses beyond the scope of their internal knowledge Cho et al. (2024); Li et al. (2024b); Zhang et al. (2024b). Despite the advancements, these models frequently suffer from hallucination Huang et al. (2023); Bai et al. (2024), undermining the faithfulness of their outputs Zhu et al. (2024). To mitigate this issue, a promising path is to have the model cite the information sources after generating a statement, offering a transparent and verifiable attribution chain.
@@
Existing studies on citation text generation mainly focus on the textual modality Gao et al. (2023); Liu et al. (2023a). However, real-world information sources are composed of multiple modalities, which convey information that is beyond the textual modality alone (e.g., flowcharts and tables). Attributing the origin of multimodal sources can improve the faithfulness and quality of model responses in broader scenarios (as shown in Figure 1), which is overlooked in previous work. Additionally, generating responses with citations to multiple modalities sources brings several challenges for MLLMs. For example, the model needs to perform cross-modal understanding, assess whether the cited evidence sufficiently supports the answers, and avoid being influenced by distracting information. This presents significant challenges to MLLMs, which are still under exploration. Therefore, we aim to build a benchmark to evaluate and analyze the multimodal citation text generation ability in MLLMs.
However, building such a benchmark is challenging. First, constructing high-quality question-answer data with multimodal sources as evidence is non-trivial. On one hand, collecting multimodal corpora requires accurately extracting the information sources. On the other hand, the evidence in the dataset must be aligned with the answers. Especially, when multiple pieces of evidence contribute to an answer, it is crucial to ensure that they are interrelated and collectively provide sufficient support for answering the question. Second, evaluating MLLMs’ performance in multimodal citation generation adds further complexity. Developing reliable evaluation methods for cross-modal entailment is the key to assessing whether cited evidence supports generated answers. Additionally, citations are required to align with the responses, ensuring the cited evidence is both necessary for answering the question and directly corresponds to the generated output. Therefore, it is essential to evaluate model performance across multiple dimensions.
@@
In this paper, we propose MCiteBench, the first benchmark for multimodal citation text generation. To tackle the above challenges, we first exploit academic papers, then accurately extract and rigorously filter the information sources from multiple modalities, forming an attribution corpus. Based on the corpus, we utilize the review-rebuttal interactions of these papers to construct QA pairs where the answer contains supportive evidence. To comprehensively assess model performance, we evaluate models from three dimensions: citation quality, source reliability, and answer accuracy. Through extensive experiments, we present several key findings: 1) While MLLMs can answer questions correctly, they struggle with citation, particularly when evidence comes from multiple sources. 2) Compared to visual information, MLLMs are better at attributing the textual evidence, indicating a potential modality bias in citation text generation.
In summary, our contributions are as follows:
  • To the best of our knowledge, we are the first benchmark for comprehensively evaluating the performance of MLLMs in multimodal citation text generation.
  • MCiteBench consists of 3,000 samples that cover a wide range of difficulty levels, including both single-source and multi-source evidence. Additionally, it encompasses cases involving single-modality and mixed-modality evidence. Given the diversity of evidence sources and complexities involved, we design multi-dimensional criteria to assess model performance.
  • We conduct experiments to analyze the performance of models in citation text generation. Results indicate that existing models struggle with citation text generation, with limitations stemming from attribution rather than understanding.
@@ 2 Related Work
Citation Text Generation
Citation text generation involves models producing text with explicit references to the sources of the information they were conditioned on. Gao et al. (2023); Liu et al. (2023a) first introduced the concept of generating responses with citations to help users quickly verify the sources of information. Subsequent works have explored two main approaches for generating responses with citations: one involves generating both the response and citations simultaneously Deng et al. (2024); Aly et al. (2024); Huang et al. (2024), while the other adds citations during the post-processing phase Slobodkin et al. (2024); Li et al. (2024a); Xia et al. (2024) to enhance the citation generation process. Furthermore, this approach has been extended to domains like long-context citation Zhang et al. (2024a) and fine-grained citation generation Xu et al. (2024). Our work focuses on addressing the attribution of outputs derived from multimodal information sources. Unlike previous approaches that focus primarily on textual data, our benchmark integrates figure and tabular evidence, aligning it more closely with real-world multimodal citation scenarios.
@@
Multimodal retrieval-augmented generation (mRAG) Zhao et al. (2023) combines retrieved multimodal information sources with multimodal large language models, which is significant for helping models answer questions that cannot be addressed solely by the model’s internal knowledge. Zhang et al. (2024b) acquires unknown visual knowledge through web search to aid in answering queries, while Li et al. (2024b) builds a self-adaptive retrieval agent to plan the reasoning path. Additionally, Cho et al. (2024) improves multi-page and multi-document understanding through multimodal retrieval. While mRAG uses retrieved passages during inference and training, it does not verify if the generated output faithfully reflects the sources. In contrast, our work focuses on whether the model can correctly attribute its generated outputs to the appropriate information sources when provided with multimodal inputs.
@@ 3 MCiteBench
In this section, we first present the definition of the multimodal citation text generation task, followed by the construction of MCiteBench. The construction pipeline of MCiteBench is illustrated in Figure 2. It consists of four steps: Attribution Corpus Collection, QA Pairs Construction, Evidence Pairing, and Quality Control. First, we gather academic papers as a rich source of multimodal content. Then, we construct QA pairs from review-rebuttal interactions and OCR-processed components, use human annotation to link answers to their supporting evidence, and apply automated filtering followed by manual verification to ensure dataset quality.
@@ 3.1 Task Definition
Given a query $q$ and a multimodal evidence set $M$, where $M$ contains both the ground truth evidence related to $q$ and distractors, the model is required to generate an answer $a$ along with a set of citations $C$. For each sentence $s_{i}$ in the answer, the model generates a set of citations $C_{i}=\{c_{i,1},c_{i,2},\dots,c_{i,k_{i}}\}$, where $k_{i}$ denotes the number of cited evidence associated with sentence $s_{i}$. Each citation $c_{i,j}$ refers to a specific piece of evidence from the multimodal evidence set $M$.
@@ 3.2 Attribution Corpus Collection
To evaluate the multimodal citation text generation of MLLMs, an attribution corpus that includes multimodal information sources and allows for easy verification of cited evidence is needed. In MCiteBench, we use academic papers as the attribution corpus because of these characteristics: 1) Academic papers contain rich content from multiple modalities (e.g., text, figure, and table) that individually or collectively support the arguments. 2) The information sources in academic papers are numbered (e.g., “Figure 1”, “Table 2”, and text in “Line 10”), making it easy to match them with the cited results. 3) Academic papers cover the latest content beyond pre-training data, reducing the risk of data leakage.
In MCiteBench, we collect academic papers from OpenReview and employ the MinerU framework Wang et al. (2024a), a state-of-the-art document extraction tool, to parse and extract multimodal content from these papers. As a result, we obtain a diverse corpus of multimodal content, including over 400k text paragraphs, 40k images, and 9k tables, which serve as candidate evidence. From this corpus, we select a subset as the evidence and distractor sources for our final 3k data samples.
@@ 3.3 QA Pairs Construction
After collecting the attribution corpus, the next step is constructing question-answer pairs with explicit references to the supporting evidence. Establishing a reliable correlation between questions and evidence is challenging, as the source of information must be accurately linked to the generated answers.
We divide MCiteBench data into two categories: Explanation and Locating. Explanation questions require in-depth analysis of evidence and often involve long-form answers, such as “How is the model’s performance evaluated?” While Locating questions are straightforward and can be answered by directly identifying the correct evidence, such as “Which model performs better on the XYZ benchmark, GPT-4o or GPT-4o-mini?”.
@@
However, generating appropriate questions from multiple sources is challenging for MLLMs. They often fail to integrate all required information, making it difficult to generate questions that fully utilize all necessary knowledge. Models often struggle to ensure that generated questions require all selected evidence, rather than being answerable by a single piece of evidence. Therefore, for Explanation questions, we leverage review-rebuttal interactions to construct QA pairs. The review and rebuttal data comprises the reviewers’ questions and authors’ responses, where the responses are consistently supported by multiple pieces of evidence from the papers (i.e., attribution corpus). From these data, we construct QA pairs $(Q,A)$ by extracting questions $q_{i}$ and the corresponding answers $a_{i}$  [^1] Details of prompt design and reference extraction strategies are in Appendix A.1.
@@ 3.4 Evidence Pairing
In review-rebuttal interactions, the authors’ response contains rich evidence to support their arguments. For example, when addressing a reviewer’s concern about model performance, an author might respond, “Our approach achieves 85.2% accuracy, as shown in Table 3 and discussed in Section 4.2.” Therefore, we extract the supportive evidence $e_{i}\in E$ from $a_{i}\in A$ to construct $(Q,A,E)$ triplets. While $E$ provides explicit references (e.g., Table 3 and Section 4.2), these references must be linked to the actual content in the source papers before they can be used as input for MLLMs. To achieve this, human annotators annotate the evidence by mapping each reference to its corresponding text, image, or table in the original papers.
@@
We construct distractors to assess the model’s ability to accurately cite relevant evidence while ignoring irrelevant information. These distractors are sampled from the same paper, ensuring a balanced distribution of multimodal content (text, images, tables).
After obtaining the distractors, we construct each data sample, which includes a question, the correct answer, multimodal ground truth evidence, and distractors. In MCiteBench, each sample is represented as $(Q,A,E,D)$, where $Q$ is the question, $A$ is the correct answer, $E$ is the evidence and $D$ is the distractors.
@@ 3.5 Quality Control
After constructing $(Q,A,E,D)$, we apply a quality control pipeline that first uses automated filtering followed by human verification.
Initially, GPT-4o assigns quality labels and filters out low-quality samples based on predefined criteria such as relevance, clarity, and evidence alignment. The filtered candidates are then manually verified by annotators to ensure consistency and accuracy, focusing on removing any unclear or incorrect instances  [^2] The human filter process is supported by a dedicated GUI interface, as shown in Appendix B.2..
@@ 3.6 Statistics of MCiteBench
As shown in Table 1, MCiteBench comprises 3,000 data samples for multimodal citation text generation tasks, extracted from 1,749 academic papers with an average of 1.72 questions per paper. Among these, 2,000 are Explanation questions that require detailed evidence analysis and often lead to long-form answers, while 1,000 are Locating questions that focus on direct evidence identification. The evidence is balanced across modalities, with 1,243 textual, 1,474 visual (including 941 figures and 533 tables), and 283 mixed-modality sources, ensuring diverse multimodal attribution scenarios.
@@ 4 Evaluation Metrics
As illustrated in Figure 3, a judge model is employed to verify whether each sentence is properly supported by its corresponding cited evidence. Citation Recall is calculated using a scoring system inspired by LongCite Zhang et al. (2024a), categorizing citations into three levels: No support, Partially supported, and Fully supported, with corresponding scores of 0, 0.5, and 1. In cases where a sentence cites multiple evidence, we concatenate the information from all cited evidence and judge the entailment relationship.
Citation Precision is determined on a binary scale, scored as either relevant (1) or irrelevant (0) to the cited evidence. For sentences citing multiple sources, the final precision score is the average across all cited evidence. Finally, Citation F1 is computed as the harmonic mean of Recall and Precision, providing a balanced measure of the model’s citation quality.
Figure 4: The calculation of Source F1 and Source Exact Match.
@@ C-F1 S-F1 S-EM Acc C-F1 S-F1 S-EM Acc C-F1 S-F1 S-EM Acc
Answer accuracy is evaluated using the LLM-As-Judge Zheng et al. (2023); Liu et al. (2023b) approach across both Explanation and Locating questions. The judge model evaluates the model’s response and the reference answer based on criteria specific to the question type. Afterward, the scores are normalized. In Explanation cases, direct comparison with a ground truth answer is not feasible. Instead, we use the authors’ responses as the reference and employ a judge model to evaluate the generated answers based on their relevance, logical consistency, and fluency. In Locating scenarios, this evaluation method mitigates issues related to errors caused by minor formatting differences  [^3] Detailed scoring criteria and judgment prompts are provided in the Appendix A.2..

### v3

# MCiteBench: A Multimodal Benchmark for Generating Text with Citations

@@ Abstract
Multimodal Large Language Models (MLLMs) have advanced in integrating diverse modalities but frequently suffer from hallucination. A promising solution to mitigate this issue is to generate text with citations, providing a transparent chain for verification. However, existing work primarily focuses on generating citations for text-only content, leaving the challenges of multimodal scenarios largely unexplored. In this paper, we introduce MCiteBench, the first benchmark designed to assess the ability of MLLMs to generate text with citations in multimodal contexts. Our benchmark comprises data derived from academic papers and review-rebuttal interactions, featuring diverse information sources and multimodal content. Experimental results reveal that MLLMs struggle to ground their outputs reliably when handling multimodal input. Further analysis uncovers a systematic modality bias and reveals how models internally rely on different sources when generating citations, offering insights into model behavior and guiding future directions for multimodal citation tasks.
@@
MCiteBench: A Multimodal Benchmark for Generating Text with Citations
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have shown remarkable progress in integrating external information from diverse modalities, allowing them to generate responses beyond the scope of their internal knowledge Cho et al. (2024); Li et al. (2024b); Zhang et al. (2024b). Despite the advancements, these models frequently suffer from hallucination Huang et al. (2023); Bai et al. (2024), undermining the faithfulness of their outputs Zhu et al. (2024). A natural strategy to alleviate this issue is citation: allowing the model to attribute each generated statement to its source, thereby improving transparency and verifiability.
@@
Existing studies on generating text with citations mainly focus on the textual modality Gao et al. (2023); Liu et al. (2023a). However, real-world information sources are inherently multimodal, often conveying information that cannot be captured by text alone. Although common in practice, citations from non-textual modalities remain underexplored. Grounding model responses in multimodal sources can improve faithfulness and quality (see Figure 1). At the same time, this task poses several challenges for MLLMs. The model must understand cross-modal content, assess the sufficiency of evidence, and remain robust to irrelevant or distracting input. These challenges are still under exploration. In this paper, we construct a benchmark to systematically evaluate MLLMs in generating text with citation from multimodal input.
However, building such a benchmark is challenging. First, constructing high-quality question-answer data with multimodal evidence is non-trivial. It requires not only the accurate extraction of heterogeneous evidence (e.g., tables, figures, and passages), but also careful alignment between the evidence and the answer. In cases where multiple pieces of evidence jointly support an answer, it is critical to ensure their mutual consistency and sufficiency. Second, evaluating MLLMs in this setting introduces additional complexity. A key issue is how to assess cross-modal entailment—whether the cited evidence truly supports the generated answer. Moreover, the citation must correspond closely to the response, ensuring that the retrieved evidence is both necessary and relevant to the output. These challenges highlight the need for a comprehensive evaluation framework that examines multiple dimensions of model performance.
@@
In this paper, we propose MCiteBench, the first benchmark for evaluating the ability of MLLMs to generate text with citations in multimodal settings. To address the challenges outlined above, we begin by collecting academic papers and extracting reliable information sources across multiple modalities. These sources are rigorously filtered to form a high-quality attribution corpus. Based on this corpus, we construct question–answer pairs using review–rebuttal interactions, where each answer is supported by evidence. To comprehensively evaluate model performance, we assess models along three axes: citation quality, source reliability, and answer accuracy. Extensive experiments reveal several notable findings: 1) While MLLMs can often answer questions correctly, they struggle to generate accurate citations, particularly when the evidence spans multiple sources. 2) MLLMs are better at attributing citations to textual than to visual evidence, suggesting a potential modality bias.
Our contributions are summarized as follows:
  • To the best of our knowledge, MCiteBench is the first benchmark that systematically evaluates the ability of MLLMs to generate text with citations from multimodal input.
  • MCiteBench comprises 3,000 samples of different difficulty levels, including both single- and multi-source evidence, as well as single- and mixed-modality cases. To support comprehensive evaluation, we define multi-dimensional metrics capturing citation quality, source reliability, and answer accuracy.
  • We conduct experiments to assess the models’ ability to generate text with citations across different modalities. Results reveal that MLLMs exhibit a modality bias, favoring textual over visual sources in citation generation.
@@ 2 Related Work
Generating Text with Citations
Recent efforts have explored the task of generating text with citations, where models are required to produce responses with explicit references to supporting sources. Gao et al. (2023); Liu et al. (2023a) first introduced this setting to improve the verifiability of model responses. Subsequent works have explored two main paradigms: generating both the response and citations simultaneously Aly et al. (2024); Huang et al. (2024), and attaching citations in a post-processing step Slobodkin et al. (2024); Li et al. (2024a). These approaches have also been extended to tasks such as long-context citation Zhang et al. (2024a) and fine-grained attribution Xu et al. (2024). Another related line of work is traditional citation text generation, which typically refers to generating citation sentences in academic papers that contain specific scientific claims and cite prior work Li and Ouyang (2024); Mandal et al. (2024); Şahinuç et al. (2024). However, existing studies focus almost exclusively on textual evidence, limiting their applicability in real-world multimodal scenarios. In this work, we address this gap by incorporating figure and tabular content as citation sources and evaluating model attribution in multimodal contexts.
@@
Multimodal retrieval-augmented generation (mRAG) Zhao et al. (2023) augments multimodal large language models with retrieved external information, enabling them to answer queries that cannot be resolved using internal knowledge alone. Zhang et al. (2024b) acquire unknown visual knowledge through web search to aid in answering queries, while Li et al. (2024b) builds a self-adaptive retrieval agent to plan the reasoning path. Additionally, Cho et al. (2024) improve multi-page and multi-document understanding through multimodal retrieval. While these approaches integrate retrieval into the generation pipeline, they do not assess whether the generated responses faithfully reflect the retrieved content. In this work, we shift the focus from retrieval itself to attribution: evaluating whether the model can correctly ground its outputs in the provided multimodal sources.
@@ 3 MCiteBench
In this section, we define the task of generating text with citations from multimodal input and describe the construction of our benchmark, MCiteBench. As shown in Figure 2, the pipeline consists of four main stages: Attribution Corpus Collection, QA Pairs Construction, Evidence Pairing, and Quality Control. We begin by collecting academic papers, which serve as a source of rich multimodal content. Based on these papers, we construct question–answer pairs from review–rebuttal interactions. Human annotators are employed to link answers to their supporting evidence.
@@ 3.1 Task Definition
Given a query $q$ and a multimodal evidence set $M$, where $M$ includes both the ground truth evidence and distractors related to $q$, the model is required to generate an answer $a$ along with a set of citations $C$. For each sentence $s_{i}$ in the answer, the model generates a set of citations $C_{i}=\{c_{i,1},c_{i,2},\dots,c_{i,k_{i}}\}$, where $k_{i}$ denotes the number of cited evidence associated with sentence $s_{i}$. Each citation $c_{i,j}$ refers to a specific piece of evidence from the multimodal evidence set $M$.
@@ 3.2 Attribution Corpus Collection
To evaluate how well MLLMs generate text with citations, an attribution corpus that includes multimodal information sources and allows for easy verification of cited evidence is needed. In MCiteBench, we use academic papers as the attribution corpus because of the following characteristics: 1) Academic papers contain rich content from multiple modalities (e.g., text, figure, and table) that individually or collectively support the arguments. 2) The information sources in academic papers are numbered (e.g., “Figure 1”, “Table 2”, and text in “Line 10”), making it easy to match them with the cited results. 3) Academic papers cover the latest contents beyond pre-training data, reducing the risk of data leakage.
We collect papers from OpenReview and extract multimodal content using MinerU Wang et al. (2024a), a state-of-the-art document parsing framework. To avoid contamination from model training data, we focus on ICLR 2025 submissions, which became publicly available in November 2024—after the knowledge cutoff of the evaluated models. ICLR is chosen for its open review process, which includes accessible reviews and author responses, offering reliable structure for citation annotation. From this collection, we obtain a diverse set of multimodal content, including over 400k text paragraphs, 40k images, and 9k tables, which serve as candidate evidence. A subset of this corpus is selected as candidate evidence and distractors for constructing the final 3k evaluation samples.
@@ 3.3 QA Pairs Construction
After collecting the attribution corpus, we construct question–answer pairs with explicit references to the supporting evidence. Establishing a reliable correlation between questions and evidence is challenging, as the source of information must be accurately linked to the generated answers.
We divide MCiteBench data into two categories: Explanation and Locating. Explanation questions require in-depth analysis of evidence and often yield long-form responses (e.g., “How is the model’s performance evaluated?”). In contrast, Locating questions are straightforward and can be answered by directly identifying the correct evidence (e.g., “Which model performs better on the XYZ benchmark, GPT-4o or GPT-4o-mini?”).
@@
However, generating questions that require information from multiple sources remains a challenge for MLLMs. Models often fail to integrate all necessary evidence, resulting in questions that can be answered by a single source rather than all selected evidence. To address this, we leverage review-rebuttal interactions to construct Explanation QA pairs. In this setting, reviewers’ questions and authors’ responses are used, with responses grounded in multiple evidence segments from the paper (i.e., attribution corpus). From these data, we construct QA pairs $(Q,A)$ by extracting questions $q_{i}$ and the corresponding answers $a_{i}$. [^1] Details of prompt design and reference extraction strategies are in Appendix A.1
@@ 3.4 Evidence Pairing
Review–rebuttal interactions often include rich evidence in the authors’ responses to support their claims. For example, when addressing a reviewer’s concern about model performance, an author might respond, “Our approach achieves 85.2% accuracy, as shown in Table 3 and discussed in Section 4.2.” These references provide valuable entry points for identifying the evidence that grounds the answer. Therefore, we extract the supportive evidence $e_{i}\in E$ from $a_{i}\in A$ to construct $(Q,A,E)$ triplets. While $E$ provides explicit references (e.g., “Table 3”, “Section 4.2”), these references must be resolved to their corresponding content in the source papers before they can be used as input for MLLMs. To achieve this, human annotators manually map each reference to the associated content in the original paper, categorizing the evidence as either text, image, or table.
@@
To evaluate whether models can correctly cite relevant sources while ignoring irrelevant ones, we introduce distractor content into the input. These distractors are sampled from the same paper, ensuring a balanced distribution of multimodal content (text, images, tables). Each final sample in MCiteBench is formatted as $(Q,A,E,D)$, where $Q$ is the question, $A$ is the correct answer, $E$ is the evidence and $D$ is the distractors.
@@ 3.5 Quality Control
After constructing $(Q,A,E,D)$, we apply a quality control pipeline that first uses automated filtering followed by human verification. Initially, GPT-4o assigns quality labels and filters out low-quality samples based on predefined criteria such as relevance, clarity, and evidence alignment. The filtered candidates are then manually verified by annotators to ensure consistency and accuracy, focusing on removing any unclear or incorrect instances. [^2] Details of the human annotation process can be found in Appendix B.2.
@@ 3.6 Statistics of MCiteBench
As shown in Table 1, MCiteBench comprises 3,000 data samples for evaluating the ability of MLLMs to generate text with citations, extracted from 1,749 academic papers with an average of 1.72 questions per paper. Among these, 2,000 are Explanation tasks that require detailed evidence analysis and often lead to long-form answers, while 1,000 are Locating tasks that focus on direct evidence identification. The evidence is balanced across modalities, with 1,243 textual, 1,474 visual (including 941 figures and 533 tables), and 283 mixed-modality sources, ensuring diverse multimodal attribution scenarios.
@@ 4 Evaluation Metrics
As illustrated in Figure 3, a judge model evaluates whether each sentence is supported by its cited evidence. Citation Recall is calculated using a scoring system inspired by LongCite Zhang et al. (2024a), categorizing citations into three levels: No support, Partially supported, and Fully supported, with corresponding scores of 0, 0.5, and 1. Citation Precision is determined on a binary scale, scored as either relevant (1) or irrelevant (0) to the cited evidence. For sentences citing multiple sources, the final precision score is the average across all cited evidence. Finally, Citation F1 is computed as the harmonic mean of Recall and Precision, providing a balanced measure of the model’s citation quality.
@@ C-F1 S-F1 S-EM Acc C-F1 S-F1 S-EM Acc C-F1 S-F1 S-EM Acc
Figure 4: The calculation of Source F1 and Source Exact Match.
@@
We evaluate answer accuracy using the LLM-As-Judge Zheng et al. (2023); Liu et al. (2023b) framework for both Explanation and Locating questions. The judge model scores each response and reference answer according to criteria specific to each question type, and the scores are then normalized. In Explanation cases, direct comparison with a ground truth answer is not feasible. Instead, we use the authors’ responses as the reference and employ a judge model to evaluate the generated answers based on their relevance, logical consistency, and fluency. In Locating scenarios, this evaluation method mitigates issues related to errors caused by minor formatting differences. [^3] Detailed scoring criteria and judgment prompts are provided in the Appendix A.2.